OpenStudy (3psilon):

Height of an Obelisk Two people decide to find the height of an obelisk. They position themselves 25 feet apart in line with, and on the same side of, the obelisk. If they find that the angles of elevation from the ground where they are standing to the top of the obelisk are 65 degrees and 44 degrees, how tall is the obelisk?

OpenStudy (anonymous):

I know you're confused bc it is 2013, but I'll forgive you.

OpenStudy (3psilon):

I'm having trouble seeing this . I though one person is x distance away and the other person is a distance (x+25) away so the entire leg of the triangle is 2x + 25 but turns out you don't add them so idk what;s going on

OpenStudy (campbell_st):

draw a diagram  you need to find the value of x using the law of sines I have included the supplementary angle for person 2 and by angle sum of the triangle the top angle is 21.. so by the law of sines \[\frac{x}{\sin(44)} = \frac{25}{\sin(21)}\] solve for x when you have x you can use the sine ratio to find the height of the obelisk hope this helps